OpenStudy (anonymous):

The top and bottom margins of a poster are 2 cm and the side margins are each 8 cm. If the area of printed material on the poster is fixed at 388 square centimeters, find the dimensions of the poster with the smallest area.

OpenStudy (anonymous):

I've done this problem about 10 times now and It still isn't right! What am I doing wrong!?!? area = (x+4)(388/x +16) darea/dx = (x+4)(-388/x²) +(388/x +16) x = sqrt(97) = 9.84 width = 9.84 length = 9.84 + 16 = 25.84

OpenStudy (anonymous):

@Mertsj (You've helped me before I was hoping maybe you could help me again! Haha :P)

OpenStudy (mertsj):

 \[(x+4)(\frac{388}{x}+16)=388+16x+\frac{1552}{x}+64=16x+1552x ^{-1}+452\]